OpenStudy (anonymous):

Consider V the space of n x n matrices, and S the subspace of simmetrical n x n matrices. Consider de dot product between A and B defined by the sum of the elements of the diagonal of AB^T. Prove that P(A)=(A+A^T)/2 is the projection of an element A of V onto S. Can anyone solve this?

OpenStudy (helder_edwin):

so u have \[\large V=\mathbb{R}^{n\times n}\qquad V\geq S=\{A\in V:A^t=A\} \] and the inner product \[\large (A|B)=\text{tr}(AB^t)=\sum_{k=1}^n(AB^t)_{kk}= \sum_{k=1}^n\sum_{h=1}^nA_{kh}B^t_{hk} \]